Please help,I hae no clue at all!
An investor invested a total of $2,100 in two mutual funds.One fund earned 9% profit while the other earned a 2% profit.If the investor's total profit was $147, how much was invested in each mutual fund?

Answer by htmentor(1343)

Let F = the initial amount invested in Fund 1
Then 2100 - F = the initial amount invested in Fund 2 [since total amount invested=2100]
Since the amount earned is equal to the profit times the fund amount,
we can write the sum of the profits as:
0.09F + 0.02(2100-F) = 147
Solve for F:
(0.09-0.02)F = 147 - 42
0.07F = 105
F = 1500
So the amount invested in Fund 1 = $1500 and the amount invested in Fund 2 = $600